OpenStudy (anonymous):

write the expression as the sine, cosine, or tangent of an angle: cos60°cos10°-sin60°sin10°

OpenStudy (anonymous):

@jim_thompson5910

jimthompson5910 (jim_thompson5910):

use the identity cos(x+y) = cos(x)cos(y) - sin(x)sin(y)

OpenStudy (anonymous):

could you do it for me?

OpenStudy (anonymous):

@jim_thompson5910

jimthompson5910 (jim_thompson5910):

compare cos60°cos10°-sin60°sin10° to cos(x+y) = cos(x)cos(y) - sin(x)sin(y)

jimthompson5910 (jim_thompson5910):

what is x? what is y?

OpenStudy (anonymous):

10 and 60?

jimthompson5910 (jim_thompson5910):

actually 60 and 10, but you got the idea cos(x+y) = cos(x)cos(y) - sin(x)sin(y) cos(60+10) = cos(60)cos(10) - sin(60)sin(10) cos(70) = cos(60)cos(10) - sin(60)sin(10)

OpenStudy (anonymous):

that's all?

jimthompson5910 (jim_thompson5910):

yep

OpenStudy (anonymous):

what about sin110°cos80°+cos110°sin80°

OpenStudy (anonymous):

@jim_thompson5910

jimthompson5910 (jim_thompson5910):

use sin(x+y) = sin(x)cos(y) + cos(x)sin(y)

OpenStudy (anonymous):

ok and whats the formula foe tan325-tan86/1+tan325tan86

jimthompson5910 (jim_thompson5910):

have a look at this page http://www.sosmath.com/trig/Trig5/trig5/trig5.html

jimthompson5910 (jim_thompson5910):

there are some useful trig identities you'll need

OpenStudy (anonymous):

thank you so much

jimthompson5910 (jim_thompson5910):

yw

OpenStudy (anonymous):

i do not see the tan formula

OpenStudy (anonymous):

@jim_thompson5910

jimthompson5910 (jim_thompson5910):

look for tan(u+v)

OpenStudy (anonymous):

ok i see